# Benchmarking Foundation Models and Parameter-Efficient Fine-Tuning for Prognosis Prediction in Medical Imaging

### v1

@@ Abstract
Artificial Intelligence (AI) holds significant promise for improving prognosis prediction in medical imaging, yet its effective application remains challenging. In this work, we introduce a structured benchmark explicitly designed to evaluate and compare the transferability of Convolutional Neural Networks and Foundation Models in predicting clinical outcomes in COVID-19 patients, leveraging diverse publicly available Chest X-ray datasets. Our experimental methodology extensively explores a wide set of fine-tuning strategies, encompassing traditional approaches such as Full Fine-Tuning and Linear Probing, as well as advanced Parameter-Efficient Fine-Tuning, methods including Low-Rank Adaptation, BitFit, VeRA, and IA3. The evaluations were conducted across multiple learning paradigms, including both extensive full-data scenarios and more clinically realistic Few-Shot Learning settings, which are critical for modeling rare disease outcomes and rapidly emerging health threats. By implementing a large-scale comparative analysis involving a diverse selection of pretrained models, including general-purpose architectures pretrained on large-scale datasets, such as CLIP and DINOv2, to biomedical-specific models like MedCLIP, BioMedCLIP, and PubMedCLIP, we rigorously assess each model’s capacity to effectively adapt and generalize to prognosis tasks, particularly under conditions of severe data scarcity and pronounced class imbalance. The benchmark was designed to capture critical conditions common in prognosis tasks, as variations in dataset size, class distribution, providing detailed insights into the strengths and limitations of each fine-tuning strategy. This extensive and structured evaluation aims to inform the practical deployment and adoption of robust, efficient, and generalizable AI-driven solutions in real-world clinical prognosis prediction workflows.
@@
journal: Medical Image Analysis
@@ 1 Introduction
Artificial Intelligence (AI) has become a transformative wave in medical imaging, widely adopted across specialties such as radiology, oncology, and pathology [1, 2]. Its effectiveness in disease detection, cancer subtype classification, and early screening [3, 4, 5] was particularly evident during the COVID-19 pandemic, which accelerated the development of AI-driven diagnostic tools for chest X-ray analysis [6, 7], highlighting AI’s value in managing large-scale health emergencies [8].
In parallel, prognostic models have seen significant advancements, making use of increasingly sophisticated methods to estimate and anticipate clinical outcomes, such as disease progression and mortality [9, 10]. However, prognosis tasks remain more complex and less explored than diagnostic tasks, primarily due to data-related challenges such as the inherent temporal structure of the data and the limited availability of high-quality annotations or sufficiently large cohorts [11, 12, 13].
@@
Historically, convolutional neural networks (CNNs) have played a central role in medical imaging, particularly for diagnostic tasks, consistently achieving strong performance when pretrained with supervised learning on large-scale natural image datasets (e.g., ImageNet [14]) and subsequently fine-tuned on clinical datasets [15]. Their architectural inductive biases and parameter efficiency made them well-suited to homogeneous datasets typical of diagnostic tasks. These same architectures have been employed also in prognostic modeling; however, they exhibit clear limitations in this context. In particular, their limited capacity to capture long-range dependencies, coupled with the requirement for full fine-tuning, reduces their suitability for tasks involving temporally structured data, complex prediction objectives, and scenarios with limited sample availability.  [16, 17].
@@
As an alternative to supervised pretraining, which requires huge labeling effort to acquire the datasets needed during training, Self-Supervised Learning (SSL) has emerged as a scalable paradigm that reduces or eliminates the need for manual annotations [18, 19]. Prominent SSL methods employed in DINO [20] or CLIP [21] exemplify this shift, leveraging unlabeled data to learn rich representations trough contrastive or distillation-based objectives. They demonstrated stronger transferability to out-of-domain tasks and robustness in high-variance clinical settings [22, 23]. This has led to the emergence of Foundation Models (FMs), which use large transformer architectures pretrained on massive unlabeled datasets. Biomedical variants of FMs have been developed to improve alignment with domain-specific tasks [24, 25, 26], yet their application in prognosis prediction remains largely untested.
The limited exploration of foundation models (FMs) in prognosis is primarily driven by the unique challenges associated with prognostic datasets, which often suffer from severe sample scarcity, extreme class imbalance, and temporally evolving clinical targets [27, 28]. Although FMs have demonstrated superior generalization and robust performance across a range of diagnostic and classification tasks, their adoption in prognostic modeling remains constrained. This is largely due to the difficulty of adapting such large-scale models to the noisy, heterogeneous, and temporally structured nature of prognostic data conditions that amplify the limitations of traditional adaptation techniques. Fine-tuning strategies remain widely adopted in this context. However, their effectiveness is often hindered by issues such as domain shift, limited supervision, and poor representation quality in small target datasets. Despite these challenges, such approaches continue to offer a practical solution due to their simplicity and demonstrated utility across a variety of clinical prediction tasks.
To overcome the scalability and data efficiency limitations of full fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) techniques [29, 30] have recently emerged as a promising alternative. Rather than updating all model weights, PEFT methods focus on modifying a small subset of task-specific parameters—such as bias terms [31], low-rank adapters [32], or scaling vectors [33], while keeping the backbone model frozen. This significantly reduces computational and memory requirements, mitigates overfitting, and enables more scalable and targeted adaptation to diverse prognostic tasks [34, 35].
Moreover, Few-Shot Learning (FSL) has emerged as a complementary paradigm within transfer learning, specifically targeting scenarios characterized by extreme data scarcity. FSL techniques aim to enable models to generalize effectively from only a handful of labeled examples [36, 37]. By leveraging prior knowledge from pretrained models and carefully designed learning strategies, such as meta-learning and prototype-based classification, FSL allows rapid adaptation to new tasks with minimal supervision. Furthermore, the integration of PEFT methods into FSL pipelines has shown to further enhance adaptation efficiency in resource-constrained environments [38], offering a practical and scalable solution for real-world medical AI applications where annotated data is often difficult and costly to obtain.
@@
In this work, we address the limited exploration of foundation models (FMs) and parameter-efficient fine-tuning (PEFT) strategies in medical prognosis. We introduce a structured benchmark to evaluate fine-tuning methods across a diverse set of pretrained models, including both general-purpose (e.g., ResNet, CLIP) and biomedical-specific architectures (e.g., MedCLIP, BioMedCLIP), covering both CNNs and FMs architectures. Our benchmark focuses on COVID-19 chest X-ray (CXR) prognosis as a clinically relevant and standardized use case, supported by public datasets with varying outcome types, sample sizes, and imbalance levels. Specifically, our contributions are as follows:
@@
    Comprehensive Comparison of FMs and CNNs architectures. We present a systematic benchmark comparing PEFT techniques against classical fine-tuning strategies, such as Full Fine Tuning (FFT) and Linear Probing (LP), across both CNNs and FMs. The benchmark covers generalist pretrained models (e.g., ResNet18) and medical pretrained models (e.g., PubMedCLIP), applied to prognostic prediction tasks. We highlight which strategies are more robust to data scarcity and class imbalance—two key challenges in prognosis modeling.
@@
    PEFT applied to FMs in Prognosis Tasks. While CNNs have been extensively studied for medical imaging, the use of FMs in prognosis prediction remains underexplored. We conduct a comparative evaluation of FMs, analyzing PEFT, LP, and FFT strategies to determine which approaches and architectures achieve the best trade-off between computational cost and predictive performance.
@@
    Few-Shot Learning in Prognostic Modeling. We provide a systematic analysis of FSL performance in clinical prognosis tasks, evaluating how well PEFT methods and FMs generalize under extreme data scarcity. This analysis simulates scenarios common in rare diseases, emerging outbreaks, and underrepresented patient populations.
By answering these questions, we aim to establish a methodological foundation for adapting FMs to clinical prognosis prediction, offering practical insights for deploying robust, efficient, and generalizable AI systems in real-world medical settings. We organized this work as follows: in Section 2, we review prior research on FMs and PEFT within medical imaging, emphasizing the current gap in studies focused on prognostic tasks. In Section 3 we introduce the benchmark datasets and our methodological framework, detailing the model selection based on pretraining paradigms, fine-tuning strategies and the experimental setup. Finally, Section 4 reports the results and corresponding discussion, structured around the three central research questions outlined earlier.
@@ 2 Related Works
To contextualize our benchmark, we review recent developments in two key areas: the application of FMs in medical imaging and the emergence of PEFT as a practical strategy for adapting these models to medical domain tasks.
2.1 Foundation Models in Medical Imaging.
FMs have become a transformative paradigm in AI, driven by large-scale SSL across multiple domains, including medical imaging [26]. Trained on extensive datasets, these models offer general-purpose representations that can be adapted to diverse downstream tasks. However, medical images differ significantly from natural images in semantics, structure, and statistical properties [24], requiring tailored pretraining and adaptation strategies for effective deployment in clinical settings. Within medical imaging, FMs’ development has primarily focused on vision-language models that leverage paired image-text supervision. CLIP [21], trained on $400$M web-sourced pairs, inspired several medical adaptations. MedCLIP [39] decouples image-text learning using semantic similarity losses to overcome clinical data limitations. BioMedCLIP [40], trained on 15M biomedical image-text pairs, extends this idea across imaging modalities, enabling more robust medical visual-language representations. Specifically for CXRs, several recent models aim to support interpretation and downstream prediction tasks. CheXagent [41], trained on a curated set of 28 public CXR datasets, combines a clinical LLM with a vision encoder to enable multi-task performance and fairness evaluation across demographic groups. DiCoM [42] introduces a novel SSL strategy for CXRs using a student-teacher framework to capture diverse visual concepts beyond single-label supervision, improving generalization and transfer. ELIXR [43] trains a lightweight adapter to align a CXR image encoder with a fixed LLM (PaLM 2 [44]), enabling tasks such as zero-shot classification, semantic search, VQA, and report quality assurance.
In 3D radiography, medical FMs have emerged to model more complex anatomical and temporal structures. Time-to-event pretraining [45] integrates longitudinal EHR data with 3D CT imaging to improve clinical risk prediction, capturing progression signals often missed by static image models. CT-FM [46], trained in contrastive learning on $148,000$ CT scan and report pairs, achieves state-of-the-art performance across triage, segmentation, retrieval, and semantic tasks, while demonstrating interpretability and test-retest robustness. In segmentation, general-domain FMs like SAM [47] have been extended to the medical domain via MedSAM [48], a ViT-based model trained on over $1.5$M medical image-text pairs. MedSAM outperforms classical architectures on various segmentation tasks, demonstrating the potential of adapting large vision-language FMs to clinical workflows. Despite these advances, the application of FMs to prognosis prediction remains limited. While models such as TITAN [49] and BEPH [50] show potential in survival prediction using histopathology, few FMs explicitly address prognosis from radiological imaging. Integrating temporally structured, multimodal clinical data with image-based representations in settings characterized by data scarcity and class imbalance remains an open challenge. In summary, while domain-specific FMs are advancing automated interpretation and general clinical AI, their adaptation to small-scale, out-of-domain prognosis tasks is still underexplored. Closing this gap requires further research into effective fine-tuning strategies, multimodal fusion, and benchmarking under realistic clinical constraints.
2.2 Parameter-Efficient Fine-Tuning in Medical Imaging.
PEFT has emerged as a practical alternative to full fine-tuning, enabling the adaptation of large-scale pretrained models to downstream tasks by updating only a small subset of parameters [51, 52, 29, 30]. By keeping the majority of the model weights frozen, PEFT strategies significantly reduce computational and memory requirements, making them particularly well-suited for resource-constrained or data-scarce settings commonly encountered in medical imaging. Broadly, PEFT methods can be categorized into three groups [30]: addition-based approaches, which inject lightweight trainable components into the network (e.g., adapters, prompt tuning); specification-based methods, which restrict updates to specific parameters such as biases, as in BitFit [31]; and reparameterization-based strategies, such as LoRA [32], which fine-tune the model in a low-rank subspace. Variants like QLoRA [35] and VeRA [53] further enhance efficiency through quantization or fixed random projections with learnable scaling. In medical imaging, PEFT has gained attention as a means to extend the capabilities of FMs to diagnostic imaging data [54, 55]. While most research to date has concentrated on diagnostic classification, there is growing interest in expanding PEFT to more complex settings, such as prognosis. For instance, FairMedFM [55] bench-marked 20 FMs across 17 datasets, uncovering performance disparities among demographic groups, while FairTune [54] proposed a PEFT parameter selection scheme optimizing both fairness and accuracy. Complementing this, Dutt et al. [51] conducted a comprehensive evaluation of $17$ PEFT strategies on CNN and transformer architectures, showing marked improvements in low-data conditions. Beyond classification, PEFT has also been employed in segmentation tasks. For example, Med-SA [56] adapted SAM [47] with domain-specific modules, and MedSAM-U [57] introduced uncertainty-aware prompt tuning for better segmentation precision. In multimodal contexts, PEFT facilitates cross-modality transfer: PEMMA [58] applied LoRA to adapt CT models to PET segmentation with minimal trainable parameters, while PeFoMed [59] demonstrated the adaptability of large vision-language models to visual question answering and report generation. PEFT methods have been applied in the clinical context to adapt large pretrained language models for clinical tasks with reduced computational demands. Gema et al. [60] introduced a two-step PEFT framework, combining specialized adapter layers for domain adaptation and downstream tasks, which improved performance across various clinical outcome prediction datasets. Moreover, FSL has emerged as a promising training paradigm that enables models to generalize from only a handful of labeled examples [36]. Specifically, Few-shot classification is a central task within FSL, and involves recognizing unseen classes given only a few labeled instances after the pretraining phase [37]. Recent findings suggest that fine-tuning pretrained ViT directly on new test classes yields competitive performance in FSC [61]. However, such fine-tuning remains computationally expensive and prone to overfitting in low-data regimes typical of medical imaging. Some of these techniques employ PEFT  [38], in order to dramatically reduces computational and memory costs, enabling the adoption of these large models in data- and resource-constrained clinical environments. Lastly, Chen et al. [62] demonstrates the potential of large language models for clinical prognosis prediction under few-shot learning conditions. Focusing on immunotherapy outcomes, the authors’ findings highlight the relevance of pretrained language models for rare disease prediction and reinforce the growing role of efficient adaptation techniques in clinical applications.
Despite growing interest in applying PEFT in medical and FSC tasks, not many works has addressed its use in prognostic modeling, especially under clinically realistic constraints such as severe class imbalance or sample size, that is critical for enabling scalable, real-world deployment of FMs in real clinical settings.

### v2

@@ Abstract
Background and Objectives
Despite the significant potential of Foundation Models (FMs) in medical imaging, their application to prognosis prediction remains challenging due to data scarcity, class imbalance, and task complexity, limiting their clinical adoption. This study introduces the first structured benchmark to assess the robustness and efficiency of transfer learning strategies for FMs compared with convolutional neural networks (CNNs) in predicting COVID-19 patient outcomes from chest X-rays. The goal is to systematically compare fine-tuning strategies, classical and parameter-efficient, under realistic clinical constraints related to data scarcity and class imbalance, offering empirical guidance for AI deployment in clinical workflows.
Methods Four publicly available COVID-19 chest X-ray datasets were used, covering mortality, severity, and ICU admission, with varying sample sizes and class imbalances. CNNs pretrained on ImageNet and FMs pretrained on general or biomedical datasets were adapted using full fine-tuning, linear probing, and parameter-efficient methods. Models were evaluated under full-data and few-shot regimes using Matthews Correlation Coefficient (MCC) and Precision–Recall AUC (PR-AUC) with cross-validation and class-weighted losses.
Results CNNs with full fine-tuning performed robustly on small, imbalanced datasets, while FMs with Parameter-Efficient Fine-Tuning (PEFT), particularly LoRA and BitFit, achieved competitive results on larger datasets. Severe class imbalance degraded PEFT performance, whereas balanced data mitigated this effect. In few-shot settings, FMs showed limited generalization, with linear probing yielding the most stable results. Conclusions: No single fine-tuning strategy proved universally optimal. CNNs remain dependable for low-resource scenarios, whereas FMs benefit from parameter-efficient methods when data are sufficient.
@@
journal: Computer Methods and Programs in Biomedicine
@@ 1 Introduction
Artificial Intelligence (AI) has become a transformative wave in medical imaging, widely adopted across specialties such as radiology, oncology, and pathology [poalelungi2023advancing, rana2023machine]. Its effectiveness in disease detection, cancer subtype classification, and early screening [ibrahim2024pneumonia, aksu2024toward, Chowdhury2020CanAH] was exceptionally evident during the COVID-19 pandemic, when the development of AI-based diagnostic models for radiology were essential to address the crisis [danilov2022automatic, signoroni2021bs, rangarajan2021artificial], underling the pivotal role of AI in managing large-scale health emergencies [rangarajan2021artificial]. In a complementary direction, prognostic models have made significant advancements, incorporating sophisticated AI-based approaches to estimate and anticipate clinical outcomes, as the progression of pathology and mortality risk [Buttia2023PrognosticMI, MultiDataset_MICCAI2024]. However, prognosis tasks remain more complex and less explored than diagnostic tasks, primarily due to data-related challenges such as the inherent temporal structure of the data and the limited availability of high-quality annotations or adequately large cohorts to effectively harness the full potential of AI-driven models [adamidi2021artificialCOVID, gillman2021automated].
@@
Among these models, Convolutional Neural Networks (CNNs) have been central to medical imaging, particularly for diagnostic tasks. They demonstrate strong performance when first trained in a supervised manner on large-scale natural image datasets (e.g., ImageNet [ImageNet]) and subsequently fine-tuned on clinical data [ke2021chextransfer]. Thanks to their inductive biases and computational efficiency, CNNs are particularly well-suited to medium-sized radiological datasets publicly available for diagnostic tasks. However, their effectiveness diminishes in prognostic settings, where datasets are typically smaller and the modeling of temporal dependencies is essential. The need for full fine-tuning, combined with their limited capacity to capture temporal structure, reduces their suitability for tasks involving complex prediction objectives or constrained sample availability [raghu2019transfusion_review, kora2022transfer].
@@
As an alternative to supervised pre-training, which requires huge labeling effort to acquire the datasets needed during training, Self-Supervised Learning (SSL) has emerged as a scalable paradigm that reduces or eliminates the need for manual annotations [bommasani2021opportunities, Azizi2021BigSM]. Models like DINO [caron2021emerging] or CLIP [radford2021learning-cliplarge] symbolize this shift, leveraging large amounts of unlabeled data to learn rich visual representations through contrastive or knowledge distillation objectives. SSL-pretrained models demonstrated remarkable transferability to unseen downstream tasks and strong robustness in high-variance clinical settings [anton2022well, wolf2023self]. Building on these advances, FMs emerged, based on the adoption of large transformer-based architectures pretrained on massive unlabeled datasets, extending the principles of self-supervised learning to a broader and more generalist perspective of AI. In healthcare, FMs have been extensively developed to align with domain-specific diagnostic and classification tasks [zhang2024challenges_imaging, moor2023foundation, khan2025comprehensiveFM]. However, the application of FMs to prognosis prediction remains largely underexplored, despite their presumed generalist adaptability and the expectation that they could transfer effectively across diverse clinical settings. The limited deployment of FMs in this domain is mainly due to the inherent challenges of adapting large-scale models to the noisy, heterogeneous, and temporally structured nature of prognostic data, conditions that exacerbate the weaknesses of conventional adaptation methods [li2024text, shi2023long].
To address these issues, more effective adaptation strategies are required. Fine-tuning remains the most widely adopted approach; however, its performance is often constrained by domain shifts, limited supervision, and suboptimal representation learning in small target datasets. Despite these drawbacks, it continues to serve as a practical baseline due to its simplicity and demonstrated effectiveness across various clinical prediction tasks. Yet, given the massive scale and complexity of FMs, the a growing need for more efficient and intelligent approaches to adapt them to new domains. Rather than updating all model parameters, PEFT methods modify only a small subset of task-specific components, such as bias terms [zaken2021bitfit], low-rank adapters [hu2021lora], or scaling vectors [liu2022-ia3], while keeping the backbone frozen. These approaches drastically reduces the computational and memory demands typically associated with full fine-tuning, mitigates overfitting, and enables more flexible adaptation across tasks with limited data availability. Despite their growing success in diverse domains, the use of PEFT techniques in prognostic modeling remains remarkably underexplored, leaving a significant gap in understanding how such methods perform in temporally structured and highly heterogeneous clinical settings.
Moving in a similar direction, Few-Shot Learning (FSL) has emerged as a complementary paradigm particularly suited to prognosis prediction, where annotated cohorts are often small and imbalanced. FSL methods aim to learn transferable representations from only a handful of labeled examples [pachetti2024systematic, woerner2024navigating], leveraging prior knowledge from pretrained models through meta-learning or prototype-based strategies. Their ability to generalize from minimal supervision is especially appealing for rare disease scenarios, where data scarcity is intrinsic. In this context, the integration of PEFT within FSL frameworks represents a compelling avenue, combining the efficiency of parameter-sparse adaptation with the data economy of few-shot learning to advance real-world prognostic modeling [Bansal2022MetaAdaptersPE].
1.1 Contributions
@@
This work fills a critical gap in the application of FMs in prognosis prediction tasks, by introducing the first large-scale benchmark of fine-tuning, encompassing Full Fine-Tuning (FFT), Linear Probing (LP), and multiple Parameter-Efficient Fine-Tuning (PEFT) methods. We systematically assess how dataset size, class imbalance, and model architecture influence adaptability, robustness, and efficiency across both CNNs and FMs. Our benchmark evaluates fine-tuning methods on a diverse set of pretrained models, including general-purpose architectures (e.g., ResNet, CLIP) and biomedical-specific ones (e.g., MedCLIP, BioMedCLIP), covering both CNN and transformer-based FMs. We focus on COVID-19 chest X-ray (CXR) prognosis as a clinically meaningful use case, leveraging public datasets with heterogeneous outcome types, sample sizes, and imbalance levels. In detail, our main contributions can be summarized as follows:
@@
    Comprehensive Benchmark of CNNs and FMs. We present the first systematic comparison of PEFT techniques against classical fine-tuning strategies (FFT and LP) across CNNs and FMs pretrained on both general and biomedical domains. The benchmark identifies which combinations of model architecture and fine-tuning method remain most resilient to data scarcity and class imbalance, two central challenges in clinical prognosis modeling.
@@
    PEFT for Prognostic Foundation Models. While CNNs have been widely explored, the application of FMs to prognosis remains largely untested. We perform an extensive evaluation of PEFT, LP, and FFT strategies on FMs, quantifying the trade-off between predictive performance, computational efficiency, and model complexity, thereby establishing the first systematic assessment of PEFT effectiveness in medical prognosis tasks.
@@
    Few-Shot Learning under Clinical Constraints. Lastly, we simulate few-shot learning scenarios to evaluate the ability of PEFT-adapted Foundation Models to generalize effectively under extreme data scarcity. This analysis mirrors real-world conditions such as rare disease modeling, emerging outbreak response, and underrepresented patient cohorts, demonstrating the practical feasibility of PEFT-based fine-tuning for rapid clinical deployment.
By addressing these questions, this study establishes a methodological foundation for adapting FMs to prognostic prediction, providing actionable insights for building robust, efficient, and generalizable AI systems in real-world medical environments. Moreover, it contributes to the broader field of radiomics by advancing the integration of deep foundation models within quantitative imaging workflows, promoting a shift from handcrafted feature engineering toward data-driven, transferable representations for prognostic analysis. The remainder of this work is organized as follows: Section 2 reviews related efforts on FMs and PEFT in medical imaging, emphasizing the lack of systematic studies on prognosis. Section 3 describes the datasets, methodological framework, model selection criteria, fine-tuning strategies, and experimental setup. Finally, Section 4 presents the results and discussion, structured around the three central research questions introduced above.
@@ 2 Related Works
To contextualize this benchmark, we review prior research on (i) the use of FMs in medical imaging and (ii) PEFT as an emerging strategy for adapting large pretrained models to medical tasks. Despite rapid progress in both fields, structured analyses of fine-tuning and PEFT for prognostic prediction remain largely absent, particularly under clinically realistic constraints such as limited data and severe class imbalance.
2.1 Foundation Models in Medical Imaging
FMs have redefined medical imaging by transferring large-scale SSL paradigms into clinical domains [khan2025comprehensiveFM]. Trained on massive unlabeled datasets, these models enable general-purpose representation learning and have led to biomedical adaptations such as MedCLIP [wang2022medclip], BioMedCLIP [zhang2023biomedclip], and CheXagent [chen2024vision], developed for radiological interpretation and multimodal reasoning. In segmentation and 3D imaging, extensions like MedSAM [ma2024segmentMEDSAM] and time-to-event pretraining [huo2024time] demonstrate the versatility of large vision-language architectures in capturing anatomical or temporal features.
However, most works focus on diagnostic and segmentation tasks, emphasizing classification or localization rather than patient-level outcome prediction. Prognostic modeling introduces additional challenges, small sample sizes, class imbalance, and temporal dependency structures, that remain poorly addressed by current FM adaptation strategies [guarrasi2025systematic]. While isolated attempts, such as BEPH [yang2025foundation], explore survival prediction, no large-scale benchmark currently evaluates FMs or PEFT approaches in prognosis-oriented tasks.
2.2 Parameter-Efficient Fine-Tuning in Medical Imaging
PEFT has emerged as a lightweight alternative to full fine-tuning, allowing large pretrained models to adapt to downstream tasks by updating only a small subset of parameters [han2024parameter, dutt2023parameter]. Approaches such as BitFit [zaken2021bitfit], LoRA [hu2021lora], and VeRA [kopiczko2023vera] have proven effective in reducing computational cost while maintaining strong predictive performance. Medical applications have largely focused on diagnostic or segmentation tasks, as demonstrated by FairMedFM [jin2024fairmedfm], FairTune [dutt2023fairtune], and Med-SA [wu2023medical], which show that PEFT can improve model fairness and generalization under low-resource conditions.
In parallel, FSL has emerged as a complementary paradigm for extreme data scarcity [pachetti2024systematic, woerner2024navigating]. Integrating PEFT into FSL frameworks further enhances efficiency and transferability, as demonstrated in MetaAdapters [Bansal2022MetaAdaptersPE].
2.3 Limitations of Related Works
While prior research has established the utility of FMs and PEFT in diagnostic and segmentation applications, several key limitations persist:
  1. 1.
    Lack of structured benchmarks assessing fine-tuning and PEFT performance under prognostic conditions;
  2. 2.
    Limited focus on data imbalance and temporal dynamics, which are intrinsic to clinical outcome prediction;
  3. 3.
    Insufficient evaluation of biomedical-specific versus general-purpose, leaving unclear which pretraining strategies best generalize to prognosis;
  4. 4.
    Absence of unified methodological frameworks that compare full fine-tuning, linear probing, and PEFT under controlled, clinically realistic settings.
Our work directly addresses these gaps by introducing the first systematic benchmark of fine-tuning strategies across CNNs and FMs for prognosis prediction, explicitly designed to analyze the interplay between dataset size, class balance, and model architecture. This framework provides a standardized foundation for evaluating FM adaptability in data-limited and high-impact medical applications.